A confidence interval for a population mean is given as (4.8, 7.2). What is the sample mean and the margin of error Group of answer choices.
Answer #1

Solution :

  = (Lower confidence interval + Upper confidence interval ) / 2

Sample mean = = (4.8+7.2) / 2=6

Margin of error = E = Upper confidence interval -   = 7.2-6=1.2